A research was conducted to improve the safety plans in a factory. In this study, the accidental data of the factory for the last 50 weeks were compiled. These data are grouped into the frequency distribution as shown below:

Number of Accidents

Numbers of Weeks

0-5

8

5-10

22

10-15

10

15-20

8

20-25

2

Draw a histogram and calculate the average number of accidents per week.


1
Expert's answer
2021-06-08T12:46:43-0400

Solution:

We need to compute the sample mean for these provided grouped data:



Now, we need to construct the midpoints based on the lower and upper limits of all the classes provided:


Now, with the midpoints, we need to multiply each midpoint for its corresponding frequency, as shown in the table below:


"\\begin{array}{ccl} \\bar X & = & \\displaystyle \\frac{1}{N} \\sum_{i=1}^k M_i \\cdot f_i \\\\\\\\ \\\\\\\\ & = & \\displaystyle \\frac{1}{50}\\left( 2.5 \\cdot 8+7.5 \\cdot 22+12.5 \\cdot 10+17.5 \\cdot 8+22.5 \\cdot 2 \\right) \\\\\\\\ \\\\\\\\ & = & \\displaystyle \\frac{495}{50} \\\\\\\\ \\\\\\\\ & = & 9.9 \\end{array}\n\u200b"

Therefore, based on the data provided, the sample mean for these grouped data is "\bar X = 9.9"


Histogram